OpenStudy (anonymous):

Help please, Find slope + put into slope int form. (6,3)(9,3) ' my answer: slope = 0/3 so slope = zero.. how do u put this in slope-int?

OpenStudy (anonymous):

y = mx + b where m = slope so, y = 0 + b use any of the two points you have to solve for b

OpenStudy (anonymous):

so answer that i would write is y = 0 + b thats all right? cuz its not asking me to solve, just put into slope int form

OpenStudy (anonymous):

well you need b to have the equation. You know y = b and if the slope is 0 y is always 3 which make b =3 so, y = 0 + 3 y = 3